OpenStudy (anonymous):

find the inverse f(x)=-1/3x-8

OpenStudy (anonymous):

y=-3(x+8)

OpenStudy (ybarrap):

y = (-1/3)x - 8 interchange x and y x = (-1/3)y - 8 solve for y y = (x + 8)/(-1/3) = -3x - 24 Inverse: y = -3x - 24 Check evaluate y at (-1/3)x - 8 y = (-3)( (-1/3) x - 8) - 24 = (x + 24) - 24 = x since inverse f evaluate at f is x, this is the inverse